### v1

# MiLQ: Benchmarking IR Models for Bilingual Web Search with Mixed Language Queries

Abstract
Despite bilingual speakers frequently using mixed-language queries in web searches, Information Retrieval (IR) research on them remains scarce. To address this, we introduce MiLQ, Mixed-Language Query test set, the first public benchmark of mixed-language queries, confirmed as realistic and highly preferred. Experiments show that multilingual IR models perform moderately on MiLQ and inconsistently across native, English, and mixed-language queries, also suggesting code-switched training data’s potential for robust IR models handling such queries. Meanwhile, intentional English mixing in queries proves an effective strategy for bilinguals searching English documents, which our analysis attributes to enhanced token matching compared to native queries.
MiLQ: Benchmarking IR Models for Bilingual Web Search
with Mixed Language Queries
Jonghwi Kim1, Deokhyung Kang1,, Seonjeong Hwang1, Yunsu Kim3, Jungseul Ok1,2, Gary Geunbae Lee1,2, 1 Graduate School of Artificial Intelligence, POSTECH, Republic of Korea , 2 Department of Computer Science and Engineering, POSTECH, Republic of Korea, 3 aiXplain Inc., Los Gatos, CA, USA, {jonghwi.kim, deokhk, seonjeongh, jungseul.ok, gblee}@postech.ac.kr, yunsu.kim@aixplain.com
1 Introduction
Code-switching [^1] In this study, code-switching, mixed-language, and code-mixing are used synonymously., where bilingual speakers alternate languages within a context, is a prevalent linguistic behavior in multilingual communities Auer (1999); Gardner-Chloros (2009); Auer (2013). This phenomenon extends to Human-Computer Interaction (HCI), especially via AI agents like ChatGPT OpenAI (2023), where understanding mixed-language input critically affects their perceived reliability by bilingual users Bawa et al. (2020); Choi et al. (2023). Information Retrieval (IR) systems also face the challenge of effectively handling such mixed-language queries Sitaram et al. (2019).
Meanwhile, recent IR research has expanded beyond Monolingual IR (MonoIR) settings to diverse multilingual settings. The benchmarks Asai et al. (2021); Lawrie et al. (2023b, a); Soboroff (2023); Adeyemi et al. (2024); Litschko et al. (2025) are widely utilized, representing diverse language scenarios. However, research on mixed-language queries remains sparse and outdated Fung et al. (1999); Gupta et al. (2014); Sequiera et al. (2015), with no publicly available benchmark.
Figure 1: Illustration of a bilingual user freely using German, English, and mixed-language queries. German elements are in green, and English in orange.
To address these gaps, we introduce MiLQ, the first Mixed-Language Query benchmark created by actual bilingual users (Figure1). Using MiLQ, we explore three main research questions: (RQ1) How realistic are our mixed-language queries, and which query language do bilingual users prefer? (RQ2) How well do existing multilingual IR models perform in Mixed-language Query Information Retrieval (MQIR)? (RQ3) Is the behavior of intentionally mixing English terms into query, noted in HCI studies Fu (2017, 2019), an effective strategy?
The main contributions of our work are:
  • We introduce MiLQ, the first public benchmark of mixed-language queries, qualified as realistic and highly preferred by bilinguals.
  • We provide a comprehensive performance analysis of multilingual IR models on MiLQ, establishing initial baselines for MQIR.
  • We show intentionally mixed-language queries are effective for English document retrieval across diverse methods, providing token-level analysis of their rationale.
Retrieval Native Num. Title Query Description Query
Scenario Lang of CMI GPT-Eval Human-Eval CMI GPT-Eval Human-Eval
[…]
SO 151 16.2$\rightarrow$59.6 2.38 2.34 2.73 2.58 2.95 5.4$\rightarrow$36.1 2.76 2.34 2.63 2.51 2.77
FI 151 7.3$\rightarrow$40.2 2.48 2.52 2.79 2.70 2.59 2.2$\rightarrow$45.3 2.63 2.15 2.63 2.44 2.28
DE 151 9.1$\rightarrow$61.8 2.67 2.68 2.61 2.50 2.21 2.1$\rightarrow$41.1 2.55 2.11 2.43 2.15 1.80
FR 151 5.7$\rightarrow$35.0 2.52 2.55 2.84 2.51 2.31 2.3$\rightarrow$32.9 2.80 2.30 2.84 2.51 2.31
[…]
FA 45 2.2$\rightarrow$15.0 2.98 2.98 2.87 2.82 2.64 0.1$\rightarrow$5.6 3.00 2.93 2.91 2.81 2.68
RU 44 0.0$\rightarrow$51.7 2.89 2.50 2.72 2.30 2.16 0.6$\rightarrow$51.7 2.93 2.45 2.73 2.16 2.14
Average 111.4 6.2$\rightarrow$39.5 2.64 2.60 2.78 2.59 2.57 5.0$\rightarrow$31.6 2.80 2.45 2.76 2.45 2.46
Table 1: Quality measurements for MiLQ (Title & Description queries). Code-Mixing Index (CMI) is on a 0-100 scale (Original Query CMI $\rightarrow$ Mixed-language Query (MiLQ) CMI). For GPT-Eval (Accuracy [Acc.] & Fluency [Flu.]) and Human-Evaluation (Acc. & Flu. & Realism [Real.]), both on a 1-3 scale, cell backgrounds are colored in a fine-grained red gradient from lightest red (scores $\approx$1.0) to darkest red (scores $\approx$3.0). The ’Average’ row is bolded. "XX" denotes the native language.
2 MiLQ: Mixed-Language Query test set
Data Construction
We started with queries from two Cross-Language IR (CLIR) benchmarks: CLEF Braschler (2003) and NeuCLIR22 Lawrie et al. (2023a), addressing native-to-English and English-to-native retrieval, respectively. These were selected to ensure diverse language scenarios while maintaining quality, based on three criteria: (1) availability of parallel English and native-language queries, (2) widespread use for performance comparison, and (3) budgetary feasibility. Both follow the TREC format Voorhees (2005), including short Title and longer Description queries, for which we created mixed-language versions.
Bilingual speakers, experienced in both languages and mixed-language search, crafted natural mixed-language queries from original English and native query pairs, while preserving the original search intent. To reflect realistic code-switching patterns, we adopt Matrix Language Frame theory Myers-Scotton (1997) and follow prior studies Fu (2017, 2019); Yong et al. (2023); Winata et al. (2023) that describe common code-switching as featuring native language as the grammar-governing matrix and English language as embedded. Accordingly, annotators integrated English terms into the native language structure only when conceptually necessary and linguistically sound. Annotation guidelines are in Appendix A.1, and MiLQ samples are in Appendix A.2 (Figures 5, 6).
Figure 2: Performance of retrieval models across CLIR, MQIR (MiLQ), and MonoIR scenarios. Results are averaged by language group: low-resource (SW, SO; MAP@100) [left], high-resource (FI, DE, FR; MAP@100) [middle], and diverse document language (ZH, FA, RU; nDCG@20) [right]. Models include BM25, specialized multi-vector dense retrievers (Mono-, Mixed-, Cross-Distill), and mContriever. See Appendix B.4 for per-language details.
Quality Measurement and Analysis
We measured MiLQ’s quality considering its language mixing, meaning preservation, naturalness, and realism (Table 1). First, for language mixing, we used the Code-Mixing Index (CMI) Das and Gambäck (2014) (0-100 scale, higher=more mixing; Appendix A.3). Average CMI increased from 6.2 to 39.5 (Title) and 5.0 to 31.6 (Description), showing substantially more mixing than originals. Next, GPT-Eval (GPT-4o) using Kuwanto et al.’s framework (high human alignment, Kendall’s Tau > 0.5) assessed MiLQ (1-3 scale; rubrics in Appendix A.4) for Accuracy (Acc.) (meaning preservation, correct term use) and Fluency (Flu.) (naturalness, readability, seamlessness). MiLQ achieved strong average GPT-Eval scores: Acc. 2.64 / Flu. 2.60 (Title) and Acc. 2.80 / Flu. 2.45 (Description). Lastly, for Human-Eval, three bilingual annotators per query assessed MiLQ on a 1-3 scale (detailed guidelines in Appendix A.5). This evaluation covered Accuracy (Acc.) and Fluency (Flu.), using criteria consistent with GPT-Eval, and an additional Realism (Real.). Realism specifically assessed how naturally bilingual speakers might use the given mixed-language query in real search scenarios. Human evaluators rated MiLQ highly, with average scores: Acc. 2.78 / Flu. 2.59 / Real. 2.57 (Title) and Acc. 2.76 / Flu. 2.45 / Real. 2.46 (Description). These consistently high scores in all metrics affirm the quality and reliability of MiLQ.
Native Title Description
Lang. Agr.(%) Nat. Mix. Eng. Agr.(%) Nat. Mix. Eng.
SW 78.8 0.44 0.82 2.01 88.7 0.30 1.05 1.68
SO 98.0 0.73 2.26 0.01 100.0 0.19 2.76 0.05
FI 70.9 1.60 1.16 1.26 81.5 1.84 0.34 1.16
DE 53.0 1.11 0.46 1.38 77.5 0.79 0.38 1.85
FR 69.5 1.02 1.86 0.47 78.8 0.82 1.89 0.38
ZH 53.2 1.40 0.92 1.08 70.2 0.88 0.94 1.48
FA 60.0 0.78 2.00 0.37 66.7 0.73 2.00 0.47
RU 72.7 0.59 1.69 0.91 75.0 0.52 1.76 0.79
AVG. 69.3 0.87 1.43 0.89 79.6 0.60 1.54 0.96
Table 2: User preference for Native (Nat.), Mixed-language (Mix.), and English (Eng.) queries. Agr.(%): Percentage of queries where a majority (2+ of 3) of annotators agreed on preferred query type(s). Nat./Mix./Eng. values represent average annotator votes (0-3) for each type. Background color intensity indicates preference strength.
To investigate user preferences for Native (Nat.), Mixed-language (Mix.), and English (Eng.) query formulations, we asked annotators to select their preferred formulation(s), allowing for multiple selections. For robust assessment, Table 2 presents results for queries in which a majority of annotators (2+ of 3) agreed on their preferred formulation. The scores for each formulation type (0-3) represent the average number of annotators who selected that type as preferred. Overall, Mix. received the highest average scores, with 1.43 for Title and 1.54 for Description queries, outperforming Nat. and Eng. formulations. However, the degree of preference varied across languages. Notably, Somali (SO) exhibited the strongest preference for mixed-language (e.g., Title: 2.26, Description: 2.76). To uncover the reasons for such variations, we conducted interviews with annotators in all languages. These discussions revealed that Somali speakers frequently code-switch, primarily using English to express modern concepts due to Somali’s limited contemporary vocabulary—findings aligned with prior literature Andrzejewski (1979, 1978); Kapchits (2019). Further interview insights, including common themes on mixed-language query usage, are provided in Appendix A.6.
In summary, this section addressed (RQ1), confirming that MiLQ is perceived as highly realistic and that bilingual users prefer mixed-language query formulations. Additional details of MiLQ are in Appendix A.7.
3 Experimental Setup
This section details our experimental setup, designed to evaluate various multilingual IR models on mixed-language queries using MiLQ.
Test Scenarios & Collections
[…]
Retrieval Models
[…]
We also include the following baselines: mContriever Izacard et al. serves as a multilingual single vector dense retriever pre-trained for broad language coverage. BM25 Robertson et al. (2009) is a standard sparse lexical matching retriever. Translate-Test first translates queries into the document’s language via Neural Machine Translation (NMT), then applies BM25 or Mono-Distill for retrieval. Detailed implementation specifics for all models are in Appendix B.3.
4 Results and Analysis
Main Results
In response to (RQ2), MiLQ (MQIR in Figure 2) shows that multilingual IR models like Mono-Distill and Cross-Distill achieve moderate performance in MQIR, performing between their MonoIR and CLIR performance. This pattern, also observed with the lexical-based BM25, is attributable to MQIR’s intermediate level of lexical cues compared to MonoIR and CLIR settings.
Further observations underscore specialization’s limitations. For instance, Mono-Distill (MonoIR-optimized) outperformed Cross-Distill (CLIR-optimized) in MonoIR settings, and vice-versa. Additionally, mContriever consistently trails specialized models. Notably, Mixed-Distill trained with artificial code-switched text shows well-balanced performance, often outperforming Cross-Distill in MonoIR and Mono-Distill in CLIR/MQIR. This highlights potential benefits of using mixed-language queries in training for a robust bilingual IR system—a core challenge MiLQ addresses: developing a single robust IR model for bilingual users freely querying in native, English or mixed language. To better harness this potential of code-switched training data explored in prior studies Litschko et al. (2023); Liu et al. (2025), future work could explore advanced methods, like multilingual LLMs, beyond simple lexicon augmentation.
[…]
Method Low Resource High Resource MonoIR
CLIR MQIR CLIR MQIR
BM25 12.35 38.35 8.56 34.92 48.71
[…]
Mono-Distill 9.53 36.34 41.32 52.26 57.93
[…]
Table 3: Performance of BM25 and Mono-Distill before and after applying NMT. The metric used is MAP@100 (%).
Effectiveness of Translate-Test
[…]
Token-Level Analysis for MQIR
Figure 3: Token-level similarity matrices from Cross-R for German and mixed-language queries on ground truth passage. The y-axis shows tokenized queries (mixed-language left, native right), and the x-axis represents the tokenized English passage. MaxSim tokens are marked by $\times$, and the code-switched parts are highlighted.
The mechanism of multi-vector retriever (e.g., ColBERT) involves identifying the most similar document tokens for each query token. While prior research Wang et al. (2023); Liu et al. (2024) has explored this in MonoIR, its behavior in other language contexts remains unexplored. This token-level analysis offers a rationale for a key aspect of (RQ3): understanding why mixed-language queries can outperform native queries for English document retrieval.
Our analysis compared MaxSim token pair similarity (a query token and its maximal similarity document token) in mixed-language versus native queries. Figure 3 (left) shows mixed-language queries, by including English terms (e.g., "Intellectual Property Rights" from German "Intellektuelle Eigentumsrechte"), allow these English tokens to form MaxSim pairings ($\times$) with accurate, higher similarity scores. Conversely, native queries (right) rely on cross-lingual interpretation of native tokens (e.g., German "Intellekt," "Eigen") to map English concepts. While MaxSim pairings are also identified ($\times$), this mapping yields weaker similarity for such crucial English concepts. Thus, intentionally mixing English terms improves MaxSim matching through higher similarity scores for English terms—a key rationale (RQ3) for MQIR’s enhanced English retrieval.
5 Conclusion
This study addressed prevalent yet understudied mixed-language querying by bilinguals by introducing MiLQ—the first public user-crafted MQIR benchmark, confirmed for realism and high user preference. Experiments on MiLQ revealed inconsistent performance of current IR models on diverse queries, underscoring the need for robust models and the potential of code-switched training data. Finally, we find that intentional English mixing boosts English document retrieval for bilinguals.
6 Limitations
While MiLQ is a valuable first public MQIR benchmark, it shares limitations common to the broader multilingual IR field. A key challenge is the test set scale; unlike large monolingual English benchmarks (e.g., MS-MARCO Bajaj et al. (2018), NQ Kwiatkowski et al. (2019) with thousands of queries), CLIR benchmarks typically comprise only tens to hundreds of queries Asai et al. (2021); Lawrie et al. (2023b, a); Soboroff (2023); Adeyemi et al. (2024). This is because creating numerous high-quality multilingual test sets is highly resource-intensive. Larger MQIR benchmarks would be beneficial, allowing for more robust methodological comparisons and fostering advancements in the field.
MiLQ currently focuses on English-native language pairs, excluding non-English/non-English combinations; future inclusion of these diverse pairings is desirable. Furthermore, while realistic, MiLQ’s user-crafted queries may not capture all code-switching patterns, as these are shaped by individual cultural and linguistic experiences. Broader participant involvement could enrich future datasets with more diverse, authentic patterns. Budgetary constraints also limited MiLQ’s initial language and domain scope, suggesting future expansions for wider utility.
These limitations and the need for larger test collections highlight promising future directions. Beyond creating larger MQIR benchmarks, key research avenues include expanding linguistic diversity (with non-English/non-English pairs), investigating broader code-switching patterns via more diverse annotators, and leveraging advanced techniques like multilingual LLMs to enhance MQIR.
Ethical Considerations
In our research, we use NeuCLIR22 (Lawrie et al., 2023a), which are licensed under CC BY-NC 4.0. The models GPT-4o (Hurst et al., 2024) are licensed under OpenAI, Apache-2.0, All models were used strictly for research purposes, and no artifacts were utilized beyond the scope of the study.
Appendix A Data Annotation
A.1 Details of the Employment and Annotation
We recruited bilingual speakers through Upwork [^2] https://www.upwork.com, who were fluent in both English and one of the following languages: Swahili (SW), Somali (SO), Finnish (FI), German (DE), French (FR), Chinese (ZH), Persian (FA), or Russian (RU). These annotators were selected based on their proficiency in both languages and their extensive experience in translation activities between English and their respective languages. We provided the annotators with clear guidelines, as shown in Figure 4. The payment was based on the number of queries, with SW, SO, FI, DE, and FR totaling 302 queries (Title + Description) for \$40. For ZH, FA, and RU, we created 94, 90, and 88 queries, respectively, with a total cost of \$20 per language.
Figure 4: Guideline for German-English mixed-language search query annotators.
A.2 Examples of Title and Description queries in MiLQ
This appendix illustrates Title and Description mixed-language queries (MiLQ) from our dataset, derived from native and English sources. The figures highlight code-switched segments and indicate their Code-Mixing Index (CMI), calculated by 1.
Figure 5: Examples of Title queries from the MiLQ dataset. Code-switched segments are highlighted, and CMI values are shown in parentheses. (*Note: Although ’Catastrophe’ is also a French word, it was identified as English by the language model in this instance.)
Figure 6: Examples of Description queries from the MiLQ dataset, corresponding to the same query IDs as the Title examples shown in Figure 5. Code-switched segments are highlighted, and CMI values are indicated in parentheses.
A.3 Code-Mixing Index (CMI)
The formula for the Code-Mixing Index (CMI) Das and Gambäck (2014) is as follows:
$CMI=\begin{cases}100\times\left(1-\frac{\max(w_{i})}{n-u}\right)&\text{if }n>u\\ 0&\text{if }n=u,\end{cases}$ (1)
where $w_{i}$ is the word count in language $i$, $\max(w_{i})$ is the word count in the primary language, $n$ is the total word count, and $u$ is the number of language-independent tokens (e.g., numbers, hashtags). In our analysis, we treat the primary language as the native language. We used GPT-4o Hurst et al. (2024) instead of existing tools for more precise language identification. While existing tools such as language-detection Nakatani (2010) and fastText Joulin et al. (2016) have been widely used for language identification, we observed certain inconsistencies in accuracy. Therefore, we leveraged LLMs for more accurate data analysis. First, we tokenize the text at the word level using NLTK [^3] https://www.nltk.org/. For Chinese text, we apply Jieba [^4] https://github.com/fxsjy/jieba, a specialized tokenizer optimized for Chinese word segmentation. After tokenization, we utilize GPT-4o to classify each token’s language using the prompt template shown in Figure 7.
Figure 7: Prompt template for language identification.
A.4 GPT-Evaluation Rubric
For GPT-based evaluation, we adopted the Accuracy and Fluency rubrics from Kuwanto et al., using their publicly available prompts and evaluation code framework. [^5] https://github.com/gkuwanto/ezswitch While their assessments utilized GPT-4O-mini, our study employed the more powerful GPT-4o. The model was instructed to evaluate generated code-switched sentences against the original monolingual sentences on a 1 (lowest) to 3 (highest) scale for each criterion.
Accuracy
This criterion measures how well the generated sentence preserves the meaning and information of the original sentence, and whether the code-switched terms are used correctly and appropriately.
Score 1 (Low):
Significant deviation from original meaning; key information missing, altered, or redundantly repeated. Code-switched terms incorrect/inappropriate. Introduces new information.
Score 2 (Moderate):
Minor deviation from original meaning; most key information present but may have slight errors. Most code-switched terms appropriate with minor mistakes.
Score 3 (High):
Fully preserves original meaning; all key information present and correct. Code-switched terms accurate and appropriately used.
Fluency
This criterion measures how natural and easy to understand the generated sentence is, considering grammar, syntax, and the smooth integration of code-switching.
Score 1 (Low):
Sentence is difficult to understand or awkward; poor grammar/syntax in either language. Code-switching disrupts sentence flow.
Score 2 (Moderate):
Sentence is understandable but may have awkward/unnatural phrasing; acceptable grammar/syntax. Code-switching somewhat smooth but not perfectly integrated.
Score 3 (High):
Sentence is natural and easy to understand; good grammar/syntax in both languages. Code-switching is smooth and seamless, enhancing flow.
A.5 Human-Evaluation Guidelines and Rubrics
For the human evaluation of mixed-language queries (MiLQ), we again recruited bilingual speakers via Upwork. Eligibility required proficiency in English and one target language at least at the B2 CEFR level, plus prior translation or linguistic experience, ensuring high-quality judgments. Annotators received detailed instructions (see Figure 8) and evaluated MiLQ quality using three criteria: Accuracy, Fluency, and Realism, rated on a 1-3 scale.
The payment scheme for this evaluation reflected task complexity and language availability: SO and SW annotators were compensated at \$20 per annotator; FI, FR, and DE annotators at \$30; and FA, ZH, and RU annotators at \$15 each.
Accuracy and Fluency Rubrics
Accuracy and Fluency rubrics mirrored those used in GPT-Evaluation (see Appendix A.4). Accuracy measures how well a MiLQ preserves the original query’s meaning and appropriately integrates code-switched terms. Fluency assesses the naturalness and clarity of language mixing, ensuring smooth integration of both languages.
Realism
This criterion, specific to human evaluation, assesses the likelihood that a bilingual speaker would naturally produce or use the given MiLQ in a real online search context.
Score 1 (Low):
Query feels unnatural or forced; unlikely to be used in real search scenarios.
Score 2 (Moderate):
Query could be used in real searches, but has noticeable awkwardness or unnatural elements.
Score 3 (High):
Query feels natural and comfortable; would likely be used in real search situations.
Figure 8: An example of the detailed annotation guidelines provided to bilingual evaluators, in this case for Somali-English mixed-language search queries. Similar guidelines were adapted for other language pairs.
A.6 Insights from Annotator Interviews on Mixed-Language Query Usage
To gain deeper insights into why and when bilingual users employ mixed-language queries (MiLQ) in real-world online searches, we conducted semi-structured interviews with all annotators. A common question posed was: "In what situations are mixed-language queries commonly used in real-world online search contexts, and for what reasons?" Table 4 summarizes the key themes derived from their responses.
Key Theme Description: Why Users Mix Languages in Queries Relevant Languages Mentioned
Lexical Gaps Native language lacks suitable or clear terms for specific concepts (esp. modern/technical), or English equivalents offer greater precision, familiarity, or avoid awkward/ambiguous translations. Swahili, Somali, French, Finnish, German, Chinese, Persian, Russian
Broader Information Access English terms are used to retrieve a broader range or larger volume of online results, especially when native-language queries are perceived as too restrictive or yield insufficient/biased information. Swahili, Somali, Finnish, German, Chinese, Persian
Querying Efficiency English terms are preferred for faster query formulation, due to shorter terms, keyboard convenience, or greater cognitive accessibility (i.e., English terms come to mind more readily or are more familiar). Swahili, Somali, French, Finnish, German, Chinese, Russian
Grammatical / Orthographic Simplification English terms are selected to bypass complex native grammatical constructions (e.g., inflections, case agreement for foreign words) or challenges posed by keyboard layouts and non-Latin scripts. Russian
Language Modernization Reliance on established English terms arises from insufficient institutional efforts to standardize native terminology for contemporary concepts, especially in digital and tech domains Somali
Table 4: Summary of Key Motivations for Mixed-Language Query Usage (Condensed to 5 Points) from Annotator Interviews
In essence, these interviews highlight that bilinguals employ MiLQ for diverse, practical reasons. Key drivers include bridging lexical gaps or seeking terminological precision when native terms are inadequate, especially for modern or technical concepts. Users also mix languages to expand information access, retrieving broader or more diverse results than native-only queries might yield, or to overcome perceived biases. Querying efficiency and fluency are other significant factors, with English often offering faster input or more readily accessible terms. Furthermore, mixed-language can serve to simplify grammatical or orthographic complexities inherent in some native languages, or address deficiencies in language modernization where native terminology for contemporary concepts is lacking.
It is important to note that the specific motivations and patterns of Mixed-language query usage are often highly speaker- and context-dependent, influenced by individual linguistic backgrounds, cognitive habits, the nature of the information need, and even momentary contextual factors. Understanding these varied drivers is crucial for developing IR systems that can effectively cater to the nuanced and dynamic search behaviors of bilingual users worldwide.
A.7 Part-of-Speech Distribution of Code-Switched Words in Queries
Figure 9: POS distribution of English code-switched words in queries from NeuCLIR22 and CLEF00-03 (left) and MiLQ dataset (right). PCW refers to punctuation-combined words.
The distribution of English words in both native and mixed-language queries predominantly shows that nouns and proper nouns are the most common parts of speech. However, in our MiLQ dataset, nouns outnumber proper nouns, which contrasts with the distribution observed in native queries. Moreover, our dataset exhibits code-switching not only in nouns and proper nouns but also in a broader range of parts of speech, including adjectives, prepositions, verbs, and pronouns, showing a more diverse pattern of code-switching compared to existing datasets.
Appendix B Experiment Details
B.1 Benchmark Statistics
NeuCLIR 22 CLEF00-03
ZH FA RU SW SO FI DE FR
# Queries 47 45 44 151 151 151 151 151
# Documents 3.2M 2.2M 4.6M 113K 113K 113K 113K 113K
# Passages 19.8M 14.0M 25.1M 1.02M 1.02M 1.02M 1.02M 1.02M
Table 5: NeuCLIR22 and CLEF00-03 benchmark statistics.
Following previous research Huang et al. (2023), we use 151 queries from the CLEF C001 – C200 topics, excluding those with no relevant judgments. English documents are sourced from the Los Angeles Times corpus, which includes 113k news articles. For high-resource languages such as Finnish, German, and French, queries are directly provided by the CLEF campaign. In contrast, for low-resource languages, Bonab et al. provided Somali and Swahili translations of English queries.
B.2 Evaluation Metrics
We evaluate retrieval performance using two standard Information Retrieval metrics: MAP@100 (Mean Average Precision at 100): Evaluates ranked lists by averaging precision scores after each relevant binary-judged document is retrieved, up to 100 results. Higher scores indicate better overall retrieval. nDCG@20 (normalized Discounted Cumulative Gain at 20): Assesses ranked lists by measuring cumulative gain from graded-relevance documents within the top 20, discounted by rank and normalized by the ideal gain. Higher scores mean better top-ranking of highly relevant items.
B.3 Implementation Details
Model Configuration
Our retrieval experiments are based on the ColBERT architecture Khattab and Zaharia (2020), a multi-vector approach for dense retrieval. We utilized the publicly available PLAID-X implementation [^6] https://github.com/hltcoe/ColBERT-X for all model training and inference. Consistent with standard ColBERT practices, most training artifacts and hyperparameters were adopted directly. Our primary modification involved setting the maximum document passage length to 180 tokens. Following established methods Bendersky and Kurland (2008); Dai and Callan (2019), documents longer than this threshold were segmented into 180-token passages. During evaluation, the score for each document was determined using the maximum passage score (MaxP) strategy.
Model Backbones and Computational Resources
We fine-tuned distinct ColBERT models for each source benchmark dataset, selecting multilingual Pre-trained Language Model (mPLM) backbones based on practices in prior relevant research Yang et al. (2024); Huang et al. (2023).
All models were trained on a system equipped with four NVIDIA A100-80GB GPUs. During the training process, model is trained with 6 passages for each query.
Hyperparameters
A common set of optimization hyperparameters was used for fine-tuning all models. We employed the AdamW optimizer with a learning rate of 5e-6. All models underwent training for 200,000 steps. The total effective batch size was 64, achieved by using a batch size of 16 per GPU across the four GPUs.
B.4 Performance in Individual Languages
Query Lang BM25 Mono-Distil Mixed-Distil Cross-Distil mContriever NMT $\rightarrow$ BM25 NMT $\rightarrow$ Mono-Distil
SW 11.56 15.93 26.64 35.56 24.52 43.70 51.00
SW&EN 35.82 38.21 48.04 47.77 32.27 47.76 56.15
EN 48.71 57.93 56.53 50.77 34.88 48.71 57.93
SO 13.14 3.12 11.07 24.91 5.58 38.43 49.28
SO&EN 40.88 34.46 43.69 49.77 25.18 48.44 57.69
EN 48.71 57.93 56.72 49.94 34.88 48.71 57.93
FI 7.02 29.65 40.50 41.99 27.23 45.06 55.14
FI&EN 40.59 45.87 49.94 49.07 32.23 45.50 56.20
EN 48.71 57.93 55.73 51.53 34.88 48.71 57.93
DE 11.70 44.84 47.16 49.69 30.15 45.70 56.60
DE&EN 42.33 56.89 55.43 54.99 32.06 47.58 57.51
EN 48.71 57.93 56.70 52.12 34.88 48.71 57.93
FR 6.95 49.46 51.96 54.95 32.23 47.27 57.02
FR&EN 21.84 54.02 55.66 55.59 33.53 48.17 56.64
EN 48.71 57.93 57.35 53.05 34.88 48.71 57.93
CLIR 10.07 28.60 35.41 41.42 23.94 44.03 53.81
MQIR 36.29 45.89 50.55 51.44 31.05 47.49 56.84
MonoIR 48.71 57.93 56.60 51.48 34.88 48.71 57.93
Table 6: Performance comparison of different retrieval models across multiple language settings for retrieving English documents. This table presents the performance of individual query languages in this scenario. Additionally, XX&EN represents queries mixing the native language and English. The metric used is MAP@100 (%). The best score(s) for each individual language query type (row) are indicated in bold. If there is a unique best score, the second best score(s) are underlined.
Query Lang BM25 Mono-Distill Mixed-Distill Cross-Distill mContriever
ZH 25.72 46.82 49.59 48.61 32.90
ZH & EN 3.67 41.13 47.46 45.48 19.54
EN 5.74 38.52 48.73 48.91 21.22
FA 34.29 48.97 46.06 47.26 15.26
FA & EN 26.02 48.69 45.36 45.39 12.97
EN 0.07 46.29 47.28 47.99 11.69
RU 36.56 47.97 48.95 49.46 36.86
RU & EN 6.14 44.99 49.71 48.02 32.74
EN 1.11 44.66 51.42 51.30 29.27
MonoIR 32.19 47.92 48.20 48.44 28.34
MQIR 11.94 44.94 47.51 46.30 21.75
CLIR 2.31 43.16 49.14 49.40 20.73
Table 7: Performance comparison of different retrieval models across multiple language settings for retrieving the native documents. This table presents the performance of individual query languages in this scenario. Additionally, XX&EN represents queries mixing the native language and English. The metric used is nDCG@20 (%). The best score(s) for each individual language query type (row) are indicated in bold. If there is a unique best score, the second best score(s) are underlined.

### v2

@@
 Despite bilingual speakers frequently using mixed-language queries in web searches, Information Retrieval (IR) research on them remains scarce.
-To address this, we introduce MiLQ, Mixed-Language Query test set, the first public benchmark of mixed-language queries, confirmed as realistic and highly preferred.
+To address this, we introduce MiLQ, Mixed-Language Query test set, the first public benchmark of mixed-language queries, qualified as realistic and relatively preferred.
 Experiments show that multilingual IR models perform moderately on MiLQ and inconsistently across native, English, and mixed-language queries, also suggesting code-switched training data’s potential for robust IR models handling such queries.
 Meanwhile, intentional English mixing in queries proves an effective strategy for bilinguals searching English documents, which our analysis attributes to enhanced token matching compared to native queries.
+[^1] The code for this work are available at : https://github.com/jonghwi-kim/milq
 MiLQ: Benchmarking IR Models for Bilingual Web Search
 with Mixed Language Queries
-Jonghwi Kim1, Deokhyung Kang1,, Seonjeong Hwang1, Yunsu Kim3, Jungseul Ok1,2, Gary Geunbae Lee1,2, 1 Graduate School of Artificial Intelligence, POSTECH, Republic of Korea , 2 Department of Computer Science and Engineering, POSTECH, Republic of Korea, 3 aiXplain Inc., Los Gatos, CA, USA, {jonghwi.kim, deokhk, seonjeongh, jungseul.ok, gblee}@postech.ac.kr, yunsu.kim@aixplain.com
+Jonghwi Kim1, Deokhyung Kang1, Seonjeong Hwang1, Yunsu Kim3thanks: This work was done when the author was at aiXplain, Jungseul Ok1,2, Gary Geunbae Lee1,2, 1 Graduate School of Artificial Intelligence, POSTECH, Republic of Korea , 2 Department of Computer Science and Engineering, POSTECH, Republic of Korea, 3 Lilt, Inc.
+{jonghwi.kim, deokhk, seonjeongh, jungseul.ok, gblee}@postech.ac.kr, yunsu.kim@lilt.com
 1 Introduction
@@
 The main contributions of our work are:
-  • We introduce MiLQ, the first public benchmark of mixed-language queries, qualified as realistic and highly preferred by bilinguals.
+  • We introduce MiLQ [^3] https://catalogue.elra.info/en-us/repository/browse/ELRA-E0047/ , the first public benchmark of mixed-language queries, qualified as realistic and relatively preferred by bilinguals.
   • We provide a comprehensive performance analysis of multilingual IR models on MiLQ, establishing initial baselines for MQIR.
@@
 […]
-SO 151 16.2$\rightarrow$59.6 2.38 2.34 2.73 2.58 2.95 5.4$\rightarrow$36.1 2.76 2.34 2.63 2.51 2.77
-FI 151 7.3$\rightarrow$40.2 2.48 2.52 2.79 2.70 2.59 2.2$\rightarrow$45.3 2.63 2.15 2.63 2.44 2.28
-DE 151 9.1$\rightarrow$61.8 2.67 2.68 2.61 2.50 2.21 2.1$\rightarrow$41.1 2.55 2.11 2.43 2.15 1.80
-FR 151 5.7$\rightarrow$35.0 2.52 2.55 2.84 2.51 2.31 2.3$\rightarrow$32.9 2.80 2.30 2.84 2.51 2.31
 […]
-FA 45 2.2$\rightarrow$15.0 2.98 2.98 2.87 2.82 2.64 0.1$\rightarrow$5.6 3.00 2.93 2.91 2.81 2.68
-RU 44 0.0$\rightarrow$51.7 2.89 2.50 2.72 2.30 2.16 0.6$\rightarrow$51.7 2.93 2.45 2.73 2.16 2.14
-Average 111.4 6.2$\rightarrow$39.5 2.64 2.60 2.78 2.59 2.57 5.0$\rightarrow$31.6 2.80 2.45 2.76 2.45 2.46
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 1: Quality measurements for MiLQ (Title & Description queries).
@@
 The ’Average’ row is bolded.
-"XX" denotes the native language.
+”XX” denotes the native language.
 2 MiLQ: Mixed-Language Query test set
@@
 Average CMI increased from 6.2 to 39.5 (Title) and 5.0 to 31.6 (Description), showing substantially more mixing than originals.
-Next, GPT-Eval (GPT-4o) using Kuwanto et al.’s framework (high human alignment, Kendall’s Tau > 0.5) assessed MiLQ (1-3 scale; rubrics in Appendix A.4) for Accuracy (Acc.) (meaning preservation, correct term use) and Fluency (Flu.) (naturalness, readability, seamlessness).
+Next, GPT-Eval (GPT-4o) using Kuwanto et al.’s framework (high human alignment, Kendall’s Tau ¿ 0.5) assessed MiLQ (1-3 scale; rubrics in Appendix A.4) for Accuracy (Acc.) (meaning preservation, correct term use) and Fluency (Flu.) (naturalness, readability, seamlessness).
 MiLQ achieved strong average GPT-Eval scores: Acc.
@@
 This section details our experimental setup, designed to evaluate various multilingual IR models on mixed-language queries using MiLQ.
-Test Scenarios & Collections
+Test Scenarios & Dataset
 […]
@@
 Token-Level Analysis for MQIR
-Figure 3: Token-level similarity matrices from Cross-R for German and mixed-language queries on ground truth passage.
+Figure 3: Token-level similarity matrices from Cross-Distill for German and mixed-language queries on ground truth passage.
 The y-axis shows tokenized queries (mixed-language left, native right), and the x-axis represents the tokenized English passage.
@@
 Our analysis compared MaxSim token pair similarity (a query token and its maximal similarity document token) in mixed-language versus native queries.
-Figure 3 (left) shows mixed-language queries, by including English terms (e.g., "Intellectual Property Rights" from German "Intellektuelle Eigentumsrechte"), allow these English tokens to form MaxSim pairings ($\times$) with accurate, higher similarity scores.
-Conversely, native queries (right) rely on cross-lingual interpretation of native tokens (e.g., German "Intellekt," "Eigen") to map English concepts.
+Figure 3 (left) shows mixed-language queries, by including English terms (e.g., ”Intellectual Property Rights” from German ”Intellektuelle Eigentumsrechte”), allow these English tokens to form MaxSim pairings ($\times$) with accurate, higher similarity scores.
+Conversely, native queries (right) rely on cross-lingual interpretation of native tokens (e.g., German ”Intellekt,” ”Eigen”) to map English concepts.
 While MaxSim pairings are also identified ($\times$), this mapping yields weaker similarity for such crucial English concepts.
@@
 5 Conclusion
-This study addressed prevalent yet understudied mixed-language querying by bilinguals by introducing MiLQ—the first public user-crafted MQIR benchmark, confirmed for realism and high user preference.
-Experiments on MiLQ revealed inconsistent performance of current IR models on diverse queries, underscoring the need for robust models and the potential of code-switched training data.
-Finally, we find that intentional English mixing boosts English document retrieval for bilinguals.
+This study addressed the prevalent yet understudied phenomenon of mixed-language querying among bilingual speakers by introducing MiLQ—the first public user-crafted MQIR benchmark, validated for both realism and high user preference.
+Our comprehensive experiments on MiLQ revealed that current IR models exhibit inconsistent performance across diverse query types, highlighting the need for more robust retrieval systems and demonstrating the promising potential of code-switched training data.
+Finally, we discovered that intentionally mixing English terms into queries serves as an effective strategy for enhancing English document retrieval among bilingual users.
 6 Limitations
@@
 Ethical Considerations
-In our research, we use NeuCLIR22 (Lawrie et al., 2023a), which are licensed under CC BY-NC 4.0.
-The models GPT-4o (Hurst et al., 2024) are licensed under OpenAI, Apache-2.0, All models were used strictly for research purposes, and no artifacts were utilized beyond the scope of the study.
+Dataset Licensing and Usage
+Our work uses three primary datasets: NeuCLIR22 Lawrie et al.
+(2023a), CLEF00-03 Braschler (2003), and our newly introduced MiLQ dataset.
+We have verified the licensing terms for all existing datasets and ensured our usage is consistent with their intended research purposes.
+The NeuCLIR22 dataset is published by NIST with public access level and is subject to the NIST Open License.
+The CLEF00-03 data is distributed by ELDA under an End-User Agreement for Evaluation Packages for Research Use, which permits evaluation purposes.
+Our MiLQ dataset will be distributed by ELDA under a free evaluation license for academic organizations.
+The dataset is accessible through our code repository.
+The MiLQ dataset and our findings are intended solely for academic research purposes in multilingual information retrieval.
+We discourage commercial deployment without further evaluation of potential societal impacts and biases.
+Human Annotation
+For our MiLQ dataset, queries were created by bilingual speakers and subsequently validated by a different group of bilingual speakers to ensure quality and reduce bias.
+All participants in both stages were compensated fairly for their work based on regional wage standards and time estimates for each task.
+The detailed annotation guidelines and compensation structure for each stage are described in Appendix A.1 and A.5.
+We obtained informed consent from all participants, clearly explaining how their contributions would be used in research and dataset creation.
+Potential Risks
+The scope of our study is limited to the news domain and nine languages: English, Swahili, Somali, Finnish, German, French, Chinese, Persian, and Russian.
+Therefore, our findings may not generalize to other domains, genres, or languages not represented in our evaluation.
+Our findings may inadvertently favor certain language pairs or retrieval approaches that work better for high-resource languages, potentially contributing to digital language divides.
+Regarding personal information, we followed the existing privacy protection measures established by NIST and ELDA for the original datasets.
+Acknowledgments
+This research was supported by the MSIT (Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program (IITP-2025-RS-2020-II201789, Contribution Rate: 47.5%) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation); by the Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2025 (Project Name: Development of an AI-Based Korean Diagnostic System for Efficient Korean Speaking Learning by Foreigners, Project Number: RS-2025-02413038, Contribution Rate: 47.5%); and by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH), Contribution Rate: 5%).
+The MiLQ dataset is released through the ELRA catalogue under the Evaluation Use - ELRA EVALUATION license, providing free access to researchers for evaluation purposes.
+If you use this dataset, please include the following citation in your acknowledgements: MiLQ: Mixed-Language Query Test Set for Bilingual Web Search - Evaluation Package, ELRA catalogue (http://catalog.elra.info), ISLRN: 317-005-302-361-6, ELRA ID: ELRA-E0047
+The following dataset was used for evaluation in this study: The CLEF Test Suite for the CLEF 2000-2003 Campaigns – Evaluation Package, ELRA catalogue (http://catalog.elra.info), ISLRN: 200-586-423-805-2, ELRA ID: ELRA-E0008
 Appendix A Data Annotation
@@
 To gain deeper insights into why and when bilingual users employ mixed-language queries (MiLQ) in real-world online searches, we conducted semi-structured interviews with all annotators.
-A common question posed was: "In what situations are mixed-language queries commonly used in real-world online search contexts, and for what reasons?" Table 4 summarizes the key themes derived from their responses.
+A common question posed was: ”In what situations are mixed-language queries commonly used in real-world online search contexts, and for what reasons?” Table 4 summarizes the key themes derived from their responses.
 Key Theme Description: Why Users Mix Languages in Queries Relevant Languages Mentioned
@@
 NeuCLIR 22 CLEF00-03
-ZH FA RU SW SO FI DE FR
-# Queries 47 45 44 151 151 151 151 151
-# Documents 3.2M 2.2M 4.6M 113K 113K 113K 113K 113K
-# Passages 19.8M 14.0M 25.1M 1.02M 1.02M 1.02M 1.02M 1.02M
+ZH FA RU SW SO FI DE FR EN
+# Queries 47 45 44 151 151 151 151 151 151
+# Documents 3.2M 2.2M 4.6M 113k
+# Passages 19.8M 14.0M 25.1M 1.01M
 Table 5: NeuCLIR22 and CLEF00-03 benchmark statistics.
@@
 Model Configuration
-Our retrieval experiments are based on the ColBERT architecture Khattab and Zaharia (2020), a multi-vector approach for dense retrieval.
+For BM25, we utilized the Pyserini toolkit Lin et al.
+(2021), which provides reproducible sparse retrieval through Lucene  [^10] https://github.com/castorini/pyserini.
+Dense retrieval experiments with mContriever Izacard et al.
+employed facebook/mcontriever-msmarco model [^11] https://huggingface.co/facebook/mcontriever-msmarco.
+Our primary retrieval experiments use the ColBERT architecture Khattab and Zaharia (2020), a multi-vector approach for dense retrieval.
 We utilized the publicly available PLAID-X implementation [^12] https://github.com/hltcoe/ColBERT-X for all model training and inference.
@@
 During evaluation, the score for each document was determined using the maximum passage score (MaxP) strategy.
+All experiments were conducted with a single run.
+Due to the approximate nearest neighbor (ANN) search employed in ColBERT, experimental results may exhibit minor variations depending on the indexing process.
+However, we observed that such variations do not lead to substantial differences.
 Model Backbones and Computational Resources
@@
 B.4 Performance in Individual Languages
-Query Lang BM25 Mono-Distil Mixed-Distil Cross-Distil mContriever NMT $\rightarrow$ BM25 NMT $\rightarrow$ Mono-Distil
+[…]
 SW 11.56 15.93 26.64 35.56 24.52 43.70 51.00